# FluencyVE: Marrying Temporal-Aware Mamba with Bypass Attention for Video Editing

### v1

@@ Abstract
Large-scale text-to-image diffusion models have achieved unprecedented success in image generation and editing. However, extending this success to video editing remains challenging. Recent video editing efforts have adapted pretrained text-to-image models by adding temporal attention mechanisms to handle video tasks. Unfortunately, these methods continue to suffer from temporal inconsistency issues and high computational overheads. In this study, we propose FluencyVE, which is a simple yet effective one-shot video editing approach. FluencyVE integrates the linear time-series module, Mamba, into a video editing model based on pretrained Stable Diffusion models, replacing the temporal attention layer. This enables global frame-level attention while reducing the computational costs. In addition, we employ low-rank approximation matrices to replace the query and key weight matrices in the causal attention, and use a weighted averaging technique during training to update the attention scores. This approach significantly preserves the generative power of the text-to-image model while effectively reducing the computational burden. Experiments and analyses demonstrate promising results in editing various attributes, subjects, and locations in real-world videos.
@@ I Introduction
Diffusion -based models (DMs) have revolutionized text-to-image (T2I) generation, showcasing unprecedented capabilities in creating high-quality and diverse images [ho2020denoising]. These models, such as Stable Diffusion (SD) [rombach2022high] and DALL·E[ramesh2021zero], leverage the power of large-scale pretrained language models to generate content that is consistent with textual cues, vastly outperforming previous generative adversarial network (GAN) based models in terms of fidelity and diversity. Recent advancements, such as ControlNet [zhang2023controlnet], are compatible with pretrained SD models and enhance them through fine-tuning of the attention layers. These methods enable users to make precise modifications to image objects, background styles, and more, using textual descriptions.
@@
Diffusion-based T2I models excel in image generation and editing but face challenges in video editing. Text-driven video editing must address three key concerns: (1) semantic alignment, ensuring that the edited video accurately reflects the intended textual prompts; (2) spatial coherence, maintaining consistency between each modified frame and its corresponding frame in the original video; and (3) temporal continuity, ensuring smooth and seamless motion across frames for a cohesive viewing experience. Currently, video generation methods can be broadly categorized into two approaches: the first involves training text-to-video (T2V) diffusion models on large-scale text-video datasets, such as Imagen Video[ho2022imagen] and MagicVideo[zhou2022magicvideo] ; the second focuses on adapting existing T2I diffusion models for video generation. Owing to the difficulty in acquiring large-scale text-video datasets and the high computational cost of training T2V models, modifying and extending T2I models has become a more feasible option. While T2I models excel at capturing spatial features, they lack the ability to model temporal dimensions, which poses challenges in maintaining motion and temporal consistency in video generation. To address the challenge of temporal modeling in T2I models, Tune-A-Video[wu2023tune] introduced a temporal attention layer, enabling efficient time-series modeling. By fine-tuning pretrained T2I models, it achieves one-shot video tuning for T2V generation, thereby eliminating the need for large-scale video datasets. Building upon this foundation, subsequent works such as CAMEL[zhang2024camel] and SAVE[karim2023save] have further refined the attention mechanism through various approaches, leading to improved video editing quality. Despite the significant success of these methods, they reduce the computational overhead by employing sparse attention, which compromises the global frame attention and can negatively impact the temporal consistency. In addition, excessive parameter fine-tuning of the pretrained T2I model can degrade its original generative performance while simultaneously increasing the computational costs.
@@
In our study, we introduced the linear time-series model Mamba[gu2023mamba] to enhance the global frame attention and temporal consistency. With a computational complexity of $O(N)$, Mamba is more efficient than self-attention, allowing deeper model architectures with efficient memory utilization. In addition, we proposed a novel fine-tuning method known as Bypass Attention. Unlike traditional efficient fine-tuning methods and the LoRA[hu2021lora] approach, our method takes advantage of low-rank approximation matrices of sizes $k\times d$ and $d\times k$ to replace the original $W_{Q}$ and $W_{K}$ matrices with the same size of $d\times d$. This adjustment significantly reduces the tunable parameters while also lowering the overall computational overhead. The contributions of this study are summarized as follows:
@@ II-A Text-to-Image Model
Advancements in GANs and diffusion models have significantly propelled text-to-image (T2I) generation [8451971]. GAN-based T2I models [reed2016generative] outperformed early methods like VQ-VAE [van2017neural] through adversarial generator-discriminator structures but faced challenges such as unstable training and mode collapse. To address these issues, DALL·E [ramesh2021zero] utilized Transformers [vaswani2017attention], showcasing the potential of attention mechanisms for complex text-image alignment. Diffusion models (DMs), such as DDPM [ho2020denoising], further improved stability and image quality through noise-learning processes. Latent Diffusion Models (LDM) [rombach2022high] optimized diffusion in latent space, reducing computational costs while enabling high-resolution tasks like LinFusion’s 16K zero-shot image generation. Stable Diffusion (SD) [rombach2022stable] enhanced text-image generation by integrating pretrained language models like CLIP [Radford2021LearningTV], enabling accurate semantics and compatibility with extensions like ControlNet [zhang2023controlnet] and LoRA [hu2021lora]. Recent innovations, such as IP-Adapter [ye2023ip] and InstantID [wang2024instantid], demonstrate efficient and controllable generation by fine-tuning attention layers in SD, solidifying roles in SD in image and video editing.
@@ II-B Video Editing Model
Early video editing models relied on heuristic approaches and handcrafted algorithms, such as optical flow-based motion tracking and color-based background segmentation [9257101, 10420468, 10345745]. Deep learning models, including CNNs and RNNs, introduced generative capabilities, with approaches like Vid2Vid [wang2018vid2vid] and sequence models such as LSTMs and C3D [tran2015learning]. Transformers, such as VideoBERT [sun2019videobert] and TimeSformer [bertasius2021timesformer], further improved editing by leveraging attention mechanisms, while multimodal models like MMVID [seo2022mmvid] extended capabilities by integrating text and audio inputs.
Diffusion models (DMs) transformed text-driven editing with fine-grained control. Techniques like Textual Inversion [gal2022image], Prompt-to-Prompt [hertz2022prompt], and Blended Diffusion [avrahami2022blended] enabled localized edits, while ControlNet [zhang2023adding] introduced external control signals. In video editing, methods like Tune-A-Video [wu2023tune] and VideoP2P [liu2024video] enhanced temporal consistency by employing cross-frame attention mechanisms and frame-guided denoising, mitigating issues such as flickering and drift. However, they still face challenges in long-range consistency and precise motion coherence.
Training-free models, including Render-A-Video [yang2023rerender], Text2Video-Zero [text2video-zero], and TokenFlow [tokenflow2023], enable efficient editing by leveraging pretrained diffusion priors, though they encounter memory constraints and limitations in temporal perception. Recent advancements further address these limitations: FLATTEN [cong2024flatten] employs optical flow-guided attention for improved frame alignment, RAVE [kara2024rave] introduces randomized noise shuffling to enhance consistency and speed, Slicedit [cohen2024slicedit] utilizes spatio-temporal slicing to maintain fine-grained edits across frames, and Factorized Diffusion Distillation [singer2024video] improves video editing efficiency via factorized learning. These techniques enhance controllability but often struggle to balance semantic coherence, temporal consistency, and editing efficiency. As a result, improvements in one aspect frequently come at the cost of another. To address these challenges, we propose a linear-complexity sequence model to better capture temporal dynamics and achieve semantically consistent video generation over long sequences.
@@ II-C State Space Models
Early sequence models like RNNs effectively captured temporal dependencies but struggled with long sequences due to vanishing gradients and lack of parallelization [sutskever2014sequence]. Transformers [vaswani2017attention] resolved these issues with self-attention, enabling parallel processing and superior long-range modeling, though their quadratic complexity posed challenges for longer sequences. This limitation spurred the development of state space models (SSMs) [gu2021combining], which offer linear complexity and efficient handling of long sequences. The introduction of S4 [gu2021efficiently] advanced SSMs by leveraging HiPPO [gu2020hippo] projections to model long-range dependencies while maintaining linear scaling. Despite its strengths, S4 lacked selective attention for specific inputs. Mamba [gu2023mamba] addressed this with selective scan algorithms and hardware-aware optimizations like parallel scanning and kernel fusion, enhancing computational efficiency and input focus. Mamba further improved flexibility for tasks requiring precise input control while preserving linear complexity. VisionMamba [liu2024vmambavisualstatespace], building on Mamba, applied bidirectional SSMs to dense visual prediction tasks, achieving superior memory efficiency and faster inference without relying on 2D priors. This allowed VisionMamba to outperform attention-based models like DeiT [touvron2021training] in object detection and segmentation [li2022exploring]. Extending VisionMamba, VideoMamba [li2024videomamba] introduced dynamic spatiotemporal modeling for video understanding, excelling in short-term action recognition and long-term video comprehension, as demonstrated on datasets like Kinetics-400 [kay2017kineticshumanactionvideo] and COIN [tang2019coin]. VideoMamba’s efficiency, combined with its capability to handle multimodal tasks like video-text retrieval, positions it as a leading solution for comprehensive video understanding.
@@ II-D Efficient Fine-Tuning
Pre-trained models trained on large datasets extract both shallow and deep features, enabling transfer learning via fine-tuning. While early approaches adjusted all parameters, the growing size of models like GPT-4 [achiam2023gpt] has made full fine-tuning impractical due to high costs and catastrophic forgetting. Selective fine-tuning methods, such as BitFit [zaken2021bitfit] and LT-SFT [ansell2021composable], update only specific parameters but struggle with complex tasks. Additive fine-tuning, including Prefix Tuning [li2021prefix] and adapter-based methods like IP-Adapter [ye2023ip], enhances task adaptation but increases computational costs. Parameter-efficient methods like LoRA [hu2021lora] use low-rank decomposition to minimize updates and allow modular task switching. However, LoRA may degrade performance in video editing tasks with large parameter requirements. Optimizing these methods is essential to balance efficiency and complexity in such scenarios.
@@ III-A Preliminaries
Our model is developed based on Stable Diffusion (SD) Model [rombach2022high]. The SD Model is a variant of DDPM[ho2020denoising], and unlike DDPM, which restores the latent representation of the data from random noise through a process of gradually adding and removing noise, SD performs denoising in the latent space. Specifically, the forward process of SD maps the input image to the latent space by compressing it through an autoencoder[van2017neural], representing it as a latent variable $z_{0}$. In this latent space, the model gradually adds noise to $z_{0}$ through a predefined Markov chain, generating a sequence of increasingly noisy latent variables $z_{t}$. Each step of the transition process follows the Gaussian distribution:
@@
where $c=\psi(\mathcal{P^{*}})$ is the embedding of the text condition $\mathcal{P^{*}}$, processed by the text-encoder CLIP ViT-H/14[Radford2021LearningTV].
@@ III-B Temporal-Aware Mamba
To address this challenge, we integrated the linear sequence processing model, Mamba [gu2023mamba], into our task. Drawing inspiration from VideoMamba [li2024videomamba], we followed the Spatial-First rule and developed four distinct scanning methods, as shown in Figure 2. If frames containing temporal and spatial information are treated as a sequence of tokens, the four scanning methods can be categorized as: (a) Spatial forward, Temporal forward; (b) Spatial forward, Temporal reverse; (c) Spatial reverse, Temporal forward; and (d) Spatial reverse, Temporal reverse.
@@ IV-A Implementation Details
In our experiments, we base our implementation on latent DMs [rombach2022high] and primarily utilize the publicly released weights from the SD model [^1] https://huggingface.co/CompVis/stable-diffusion-v1-4. From the input video, we extracted 32 uniformly distributed frames at a resolution of 512 × 512, and applied our method to fine-tune the models over 500 steps with a learning rate of $3\times 10^{-5}$ and a batch size of 1. During inference, we employed the DDIM sampler [song2020denoising] with 50 steps, alongside classifier-free guidance [ho2022classifier] with a guidance scale of 12.5. All experiments were conducted on a single NVIDIA A800 GPU using PyTorch 2.0.0.
Dataset. We conducted comparative experiments on 53 videos from the LOVEU-TGVE competition[wu2023cvpr2023textguided], comprising 16 videos from the DAVIS dataset[perazzi2016benchmark] and 37 from Videvo. Each video was uniformly sampled into 32 frames at a resolution of 480 × 480. In addition, each video was paired with a ground-truth caption and four creative text prompts designed for tasks such as object manipulation, background modification, style transformation, and multiple changes.
@@ IV-C Comparisons
Baselines. We evaluated our method against four state-of-the-art(SOTA) video editing approaches: (1)Tune-A-Video (TAV) [wu2023tune]: A widely recognized SOTA method in video editing, serving as a conventional baseline for related works. (2)CAMEL [zhang2024camel]: An extension of TAV that introduces causal motion enhancement. (3)SAVE [karim2023save]: Builds upon TAV by incorporating spectral-shift-aware adaptation. (4)VidToMe [li2024vidtome]: A zero-shot video editing method based on self-attention tokens merging. (5)Slicedit [cohen2024slicedit]: A zero-shot video editing method using Spatio-Temporal Slices.
@@
Qualitative Results. The results of different methods on the editing subjects are shown in Fig. 6. It can be observed that TAV [wu2023tune] achieved relatively better temporal consistency through its temporal modeling and one-shot tuning. However, the poses and motions in the edited video were not faithful to the original video. In addition, owing to excessive parameter adjustments, some frames became blurred, which compromised part of the performance of the SD model. While VidToMe[li2024vidtome] and Slicedit[cohen2024slicedit], as zero-shot methods, remain faithful to the original video and ensure temporal consistency, their editing results fail to align with the intended semantics. For instance, VidToMe replaces the trackside marker with old tires suitable for an off-road setting, yet retains the number three, which breaks semantic coherence. Slicedit merely adjusts the video tone without altering key scene elements, resulting in semantically irrelevant edits. SAVE [karim2023save] improved the fine-tuning of the SD model to ensure frame quality, but it did not enhance the temporal consistency, leading to flickering and disappearing frames between shots. CAMEL [zhang2024camel] strengthened attention to motion, but some frames still failed to maintain temporal consistency. For example, certain frames in the muddy off-road track exhibit unnaturally bright regions lacking road textures. By contrast, our method enhances global frame attention, ensuring that the generated video is not only faithful to the original but also exhibits a high degree of temporal consistency and motion continuity across frames. As shown in Fig. 6, our model successfully completes the transformation from car to truck, while accurately preserving the vehicle’s original orientation and generating a background that best matches the intended semantics.
@@
Quantitative Results. We evaluated our method in comparison to baseline models using both automatic metrics and a user study, with the results for the frame consistency and textual faithfulness presented in Tab. I. Automatic Metrics. We note that there is no universally accepted evaluation standard for video editing. Therefore, we opted to use the three metrics provided by the LOVEU-TGVE competition[wu2023cvpr2023textguided] as our evaluation criteria: (i) CLIP Score text: This metric measures the alignment between video frames and a text prompt using the average cosine similarity of their embeddings obtained from a pretrained CLIP ViT-L/14 model[Radford2021LearningTV]. (ii) CLIP Score frame: Evaluates the frame consistency among video frames by computing the average cosine similarity between frame embeddings, ignoring self-similarity, to gauge the internal coherence of the video[hessel2022clipscorereferencefreeevaluationmetric]. (iii) Pick Score: This metric leverages a CLIP model trained on human preferences to evaluate the alignment between video frames and a given prompt from a human perspective[kirstain2023pickapicopendatasetuser]. User Study Aligned with the two key objectives of video editing, we asked 150 users to select: a) the video with higher editing fidelity, and b) the one with better temporal consistency. Each user evaluated 30 randomly selected videos, and the final score was based on the percentage of users preferring each method.

### v2

@@ Abstract
Large-scale text-to-image diffusion models have achieved unprecedented success in image generation and editing. However, extending this success to video editing remains challenging. Recent video editing efforts have adapted pretrained text-to-image models by adding temporal attention mechanisms to handle video tasks. Unfortunately, these methods continue to suffer from temporal inconsistency issues and high computational overheads. In this study, we propose FluencyVE, which is a simple yet effective one-shot video editing approach. FluencyVE integrates the linear time-series module, Mamba, into a video editing model based on pretrained Stable Diffusion models, replacing the temporal attention layer. This enables global frame-level attention while reducing the computational costs. In addition, we employ low-rank approximation matrices to replace the query and key weight matrices in the causal attention, and use a weighted averaging technique during training to update the attention scores. This approach significantly preserves the generative power of the text-to-image model while effectively reducing the computational burden. Experiments and analyses demonstrate promising results in editing various attributes, subjects, and locations in real-world videos. Our code is available at https://github.com/CIMASA/FluencyVE.
@@ I Introduction
Diffusion -based models (DMs) have revolutionized text-to-image (T2I) generation, showcasing unprecedented capabilities in creating high-quality and diverse images [16]. These models, such as Stable Diffusion (SD) [35] and DALL·E[33], leverage the power of large-scale pretrained language models to generate content that is consistent with textual cues, vastly outperforming previous generative adversarial network (GAN) based models in terms of fidelity and diversity. Recent advancements, such as ControlNet [57], are compatible with pretrained SD models and enhance them through fine-tuning of the attention layers. These methods enable users to make precise modifications to image objects, background styles, and more, using textual descriptions.
@@
Diffusion-based T2I models excel in image generation and editing but face challenges in video editing. Text-driven video editing must address three key concerns: (1) semantic alignment, ensuring that the edited video accurately reflects the intended textual prompts; (2) spatial coherence, maintaining consistency between each modified frame and its corresponding frame in the original video; and (3) temporal continuity, ensuring smooth and seamless motion across frames for a cohesive viewing experience. Currently, video generation methods can be broadly categorized into two approaches: the first involves training text-to-video (T2V) diffusion models on large-scale text-video datasets, such as Imagen Video[15] and MagicVideo[60] ; the second focuses on adapting existing T2I diffusion models for video generation. Owing to the difficulty in acquiring large-scale text-video datasets and the high computational cost of training T2V models, modifying and extending T2I models has become a more feasible option. While T2I models excel at capturing spatial features, they lack the ability to model temporal dimensions, which poses challenges in maintaining motion and temporal consistency in video generation. To address the challenge of temporal modeling in T2I models, Tune-A-Video[51] introduced a temporal attention layer, enabling efficient time-series modeling. By fine-tuning pretrained T2I models, it achieves one-shot video tuning for T2V generation, thereby eliminating the need for large-scale video datasets. Building upon this foundation, subsequent works such as CAMEL[56] and SAVE[20] have further refined the attention mechanism through various approaches, leading to improved video editing quality. Despite the significant success of these methods, they reduce the computational overhead by employing sparse attention, which compromises the global frame attention and can negatively impact the temporal consistency. In addition, excessive parameter fine-tuning of the pretrained T2I model can degrade its original generative performance while simultaneously increasing the computational costs.
@@
In our study, we introduced the linear time-series model Mamba[10] to enhance the global frame attention and temporal consistency. With a computational complexity of $O(N)$, Mamba is more efficient than self-attention, allowing deeper model architectures with efficient memory utilization. In addition, we proposed a novel fine-tuning method known as Bypass Attention. Unlike traditional efficient fine-tuning methods and the LoRA[18] approach, our method takes advantage of low-rank approximation matrices of sizes $k\times d$ and $d\times k$ to replace the original $W_{Q}$ and $W_{K}$ matrices with the same size of $d\times d$. This adjustment significantly reduces the tunable parameters while also lowering the overall computational overhead. The contributions of this study are summarized as follows:
@@ II-A Text-to-Image Model
Advancements in GANs and diffusion models have significantly propelled text-to-image (T2I) generation [50]. GAN-based T2I models [34] outperformed early methods like VQ-VAE [45] through adversarial generator-discriminator structures but faced challenges such as unstable training and mode collapse. To address these issues, DALL·E [33] utilized Transformers [46], showcasing the potential of attention mechanisms for complex text-image alignment. Diffusion models (DMs), such as DDPM [16], further improved stability and image quality through noise-learning processes. Latent Diffusion Models (LDM) [35] optimized diffusion in latent space, reducing computational costs while enabling high-resolution tasks like LinFusion’s 16K zero-shot image generation. Stable Diffusion (SD) [36] enhanced text-image generation by integrating pretrained language models like CLIP [32], enabling accurate semantics and compatibility with extensions like ControlNet [57] and LoRA [18]. Recent innovations, such as IP-Adapter [54] and InstantID [47], demonstrate efficient and controllable generation by fine-tuning attention layers in SD, solidifying roles in SD in image and video editing.
@@ II-B Video Editing Model
Early video editing models relied on heuristic approaches and handcrafted algorithms, such as optical flow-based motion tracking and color-based background segmentation [49, 59, 25]. Deep learning models, including CNNs and RNNs, introduced generative capabilities, with approaches like Vid2Vid [48] and sequence models such as LSTMs and C3D [44]. Transformers, such as VideoBERT [40] and TimeSformer [4], further improved editing by leveraging attention mechanisms, while multimodal models like MMVID [37] extended capabilities by integrating text and audio inputs.
Diffusion models (DMs) transformed text-driven editing with fine-grained control. Techniques like Textual Inversion [7], Prompt-to-Prompt [13], and Blended Diffusion [3] enabled localized edits, while ControlNet [58] introduced external control signals. In video editing, methods like Tune-A-Video [51] and VideoP2P [29] enhanced temporal consistency by employing cross-frame attention mechanisms and frame-guided denoising, mitigating issues such as flickering and drift. However, they still face challenges in long-range consistency and precise motion coherence.
Training-free models, including Render-A-Video [53], Text2Video-Zero [22], and TokenFlow [8], enable efficient editing by leveraging pretrained diffusion priors, though they encounter memory constraints and limitations in temporal perception. Recent advancements further address these limitations: FLATTEN [6] employs optical flow-guided attention for improved frame alignment, RAVE [19] introduces randomized noise shuffling to enhance consistency and speed, Slicedit [5] utilizes spatio-temporal slicing to maintain fine-grained edits across frames, and Factorized Diffusion Distillation [38] improves video editing efficiency via factorized learning. These techniques enhance controllability but often struggle to balance semantic coherence, temporal consistency, and editing efficiency. As a result, improvements in one aspect frequently come at the cost of another. To address these challenges, we propose a linear-complexity sequence model to better capture temporal dynamics and achieve semantically consistent video generation over long sequences.
@@ II-C State Space Models
Early sequence models like RNNs effectively captured temporal dependencies but struggled with long sequences due to vanishing gradients and lack of parallelization [41]. Transformers [46] resolved these issues with self-attention, enabling parallel processing and superior long-range modeling, though their quadratic complexity posed challenges for longer sequences. This limitation spurred the development of state space models (SSMs) [12], which offer linear complexity and efficient handling of long sequences. The introduction of S4 [11] advanced SSMs by leveraging HiPPO [9] projections to model long-range dependencies while maintaining linear scaling. Despite its strengths, S4 lacked selective attention for specific inputs. Mamba [10] addressed this with selective scan algorithms and hardware-aware optimizations like parallel scanning and kernel fusion, enhancing computational efficiency and input focus. Mamba further improved flexibility for tasks requiring precise input control while preserving linear complexity. VisionMamba [30], building on Mamba, applied bidirectional SSMs to dense visual prediction tasks, achieving superior memory efficiency and faster inference without relying on 2D priors. This allowed VisionMamba to outperform attention-based models like DeiT [43] in object detection and segmentation [28]. Extending VisionMamba, VideoMamba [24] introduced dynamic spatiotemporal modeling for video understanding, excelling in short-term action recognition and long-term video comprehension, as demonstrated on datasets like Kinetics-400 [21] and COIN [42]. VideoMamba’s efficiency, combined with its capability to handle multimodal tasks like video-text retrieval, positions it as a leading solution for comprehensive video understanding.
@@ II-D Efficient Fine-Tuning
Pre-trained models trained on large datasets extract both shallow and deep features, enabling transfer learning via fine-tuning. While early approaches adjusted all parameters, the growing size of models like GPT-4 [1] has made full fine-tuning impractical due to high costs and catastrophic forgetting. Selective fine-tuning methods, such as BitFit [55] and LT-SFT [2], update only specific parameters but struggle with complex tasks. Additive fine-tuning, including Prefix Tuning [26] and adapter-based methods like IP-Adapter [54], enhances task adaptation but increases computational costs. Parameter-efficient methods like LoRA [18] use low-rank decomposition to minimize updates and allow modular task switching. However, LoRA may degrade performance in video editing tasks with large parameter requirements. Optimizing these methods is essential to balance efficiency and complexity in such scenarios.
@@ III-A Preliminaries
Our model is developed based on Stable Diffusion (SD) Model [35]. The SD Model is a variant of DDPM[16], and unlike DDPM, which restores the latent representation of the data from random noise through a process of gradually adding and removing noise, SD performs denoising in the latent space. Specifically, the forward process of SD maps the input image to the latent space by compressing it through an autoencoder[45], representing it as a latent variable $z_{0}$. In this latent space, the model gradually adds noise to $z_{0}$ through a predefined Markov chain, generating a sequence of increasingly noisy latent variables $z_{t}$. Each step of the transition process follows the Gaussian distribution:
@@
where $c=\psi(\mathcal{P^{*}})$ is the embedding of the text condition $\mathcal{P^{*}}$, processed by the text-encoder CLIP ViT-H/14[32].
@@ III-B Temporal-Aware Mamba
To address this challenge, we integrated the linear sequence processing model, Mamba [10], into our task. Drawing inspiration from VideoMamba [24], we followed the Spatial-First rule and developed four distinct scanning methods, as shown in Figure 2. If frames containing temporal and spatial information are treated as a sequence of tokens, the four scanning methods can be categorized as: (a) Spatial forward, Temporal forward; (b) Spatial forward, Temporal reverse; (c) Spatial reverse, Temporal forward; and (d) Spatial reverse, Temporal reverse.
@@ IV-A Implementation Details
In our experiments, we base our implementation on latent DMs [35] and primarily utilize the publicly released weights from the SD model [^1] https://huggingface.co/CompVis/stable-diffusion-v1-4. From the input video, we extracted 32 uniformly distributed frames at a resolution of 512 × 512, and applied our method to fine-tune the models over 500 steps with a learning rate of $3\times 10^{-5}$ and a batch size of 1. During inference, we employed the DDIM sampler [39] with 50 steps, alongside classifier-free guidance [17] with a guidance scale of 12.5. All experiments were conducted on a single NVIDIA A800 GPU using PyTorch 2.0.0.
Dataset. We conducted comparative experiments on 53 videos from the LOVEU-TGVE competition[52], comprising 16 videos from the DAVIS dataset[31] and 37 from Videvo. Each video was uniformly sampled into 32 frames at a resolution of 480 × 480. In addition, each video was paired with a ground-truth caption and four creative text prompts designed for tasks such as object manipulation, background modification, style transformation, and multiple changes.
@@ IV-C Comparisons
Baselines. We evaluated our method against four state-of-the-art(SOTA) video editing approaches: (1)Tune-A-Video (TAV) [51]: A widely recognized SOTA method in video editing, serving as a conventional baseline for related works. (2)CAMEL [56]: An extension of TAV that introduces causal motion enhancement. (3)SAVE [20]: Builds upon TAV by incorporating spectral-shift-aware adaptation. (4)VidToMe [27]: A zero-shot video editing method based on self-attention tokens merging. (5)Slicedit [5]: A zero-shot video editing method using Spatio-Temporal Slices.
@@
Qualitative Results. The results of different methods on the editing subjects are shown in Fig. 6. It can be observed that TAV [51] achieved relatively better temporal consistency through its temporal modeling and one-shot tuning. However, the poses and motions in the edited video were not faithful to the original video. In addition, owing to excessive parameter adjustments, some frames became blurred, which compromised part of the performance of the SD model. While VidToMe[27] and Slicedit[5], as zero-shot methods, remain faithful to the original video and ensure temporal consistency, their editing results fail to align with the intended semantics. For instance, VidToMe replaces the trackside marker with old tires suitable for an off-road setting, yet retains the number three, which breaks semantic coherence. Slicedit merely adjusts the video tone without altering key scene elements, resulting in semantically irrelevant edits. SAVE [20] improved the fine-tuning of the SD model to ensure frame quality, but it did not enhance the temporal consistency, leading to flickering and disappearing frames between shots. CAMEL [56] strengthened attention to motion, but some frames still failed to maintain temporal consistency. For example, certain frames in the muddy off-road track exhibit unnaturally bright regions lacking road textures. By contrast, our method enhances global frame attention, ensuring that the generated video is not only faithful to the original but also exhibits a high degree of temporal consistency and motion continuity across frames. As shown in Fig. 6, our model successfully completes the transformation from car to truck, while accurately preserving the vehicle’s original orientation and generating a background that best matches the intended semantics.
@@
Quantitative Results. We evaluated our method in comparison to baseline models using both automatic metrics and a user study, with the results for the frame consistency and textual faithfulness presented in Tab. I. Automatic Metrics. We note that there is no universally accepted evaluation standard for video editing. Therefore, we opted to use the three metrics provided by the LOVEU-TGVE competition[52] as our evaluation criteria: (i) CLIP Score text: This metric measures the alignment between video frames and a text prompt using the average cosine similarity of their embeddings obtained from a pretrained CLIP ViT-L/14 model[32]. (ii) CLIP Score frame: Evaluates the frame consistency among video frames by computing the average cosine similarity between frame embeddings, ignoring self-similarity, to gauge the internal coherence of the video[14]. (iii) Pick Score: This metric leverages a CLIP model trained on human preferences to evaluate the alignment between video frames and a given prompt from a human perspective[23]. User Study Aligned with the two key objectives of video editing, we asked 150 users to select: a) the video with higher editing fidelity, and b) the one with better temporal consistency. Each user evaluated 30 randomly selected videos, and the final score was based on the percentage of users preferring each method.